# Near Coincidences and Nilpotent Division Fields

### v1

@@ Abstract.
Let $E/\mathbb{Q}$ be an elliptic curve. We say that $E$ has a near coincidence of level $(n,m)$ if $m\mid n$ and $\mathbb{Q}(E[n])=\mathbb{Q}(E[m],\zeta_{n})$. In the present paper we classify near coincidences of prime power level. We use this result to give a classification of values of $n$ for which $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ is a nilpotent group. In particular, if we assume that there are no non-CM rational points on the modular curves $X_{ns}^{+}(p)$ for primes $p>11$, then $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ nilpotent implies that $n$ is a power of $2$ or $n\in\{3,5,6,7,15,21\}$.
@@ 1. Introduction
We start as many did before us, by fixing an algebraic closure of $\mathbb{Q}$, $\overline{\mathbb{Q}}$. A classical result in arithmetic geometry is that given an elliptic curve $E/\mathbb{Q}$ and a field $K$ containing $\mathbb{Q}$, the $K$-rational points on $E$ form an abelian group. For $n\in\mathbb{Z}^{+}$ we let
$E[n]:=\{P\in E(\overline{\mathbb{Q}})\mid nP=\mathcal{O}\},$
and $\mathbb{Q}(E[n])/\mathbb{Q}$ be the field extension that one gets by adjoining the coordinates of the elements of $E[n]$. The field $\mathbb{Q}(E[n])$ is called the $n$-division field of $E$. A natural question one might ask is if there are elliptic curves $E/\mathbb{Q}$ and distinct $m,n\in\mathbb{Z}^{+}$ such that $\mathbb{Q}(E[m])=\mathbb{Q}(E[n])$? This is a question that was explored in [9] and more recently by Yvon in [35]. If $E$ is an elliptic curves and $m,n\in\mathbb{Z}^{+}$ such that $\mathbb{Q}(E[m])=\mathbb{Q}(E[n])$, we say that $E$ has a coincidence of divisions fields. A partial answer to the question asked is given in the following theorem.
@@
[9, Theorem 1.4] Let $E/\mathbb{Q}$ be an elliptic curve, $p$ be a prime, and let $n\in\mathbb{Z}^{+}.$
  1. (1)
    Suppose $\mathbb{Q}(E[p^{n+1}])=\mathbb{Q}(E[p^{n}]).$ Then $p=2$ and $n=1.$
  2. (2)
A large part of the proof of part (1) of this theorem is showing that $\mathbb{Q}(\zeta_{p^{n+1}})$ cannot be contained inside of $\mathbb{Q}(E[p^{n}])$ when $p$ is odd. From the Weil pairing, we know that $\mathbb{Q}(\zeta_{p^{n+1}})\subseteq\mathbb{Q}(E[p^{n+1}])$. These two things together ensure that $\mathbb{Q}(E[p^{n}])\neq\mathbb{Q}(E[p^{n+1}])$. This also explains the existence of part (2) of this theorem.
Seeing that it is the necessary roots of unity that prevent equality, one could ask a new question about near coincedences of division fields.
Question 1.2.
Are there any elliptic curves $E/\mathbb{Q}$ and distinct positive integers $m$ and $n$ such that $m\mid n$ and $\mathbb{Q}(E[n])=\mathbb{Q}(E[m],\zeta_{n})?$
Our first main theorem is a first step to answering this question.
@@
$\mathbb{Q}(E[p^{n+1}])=\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}).$
Then, it must be that $p=2$ or $3$ and $n=1$. Further, if $p=2$, then $E$ must correspond to a rational point on the one of the modular curves with RSZB label 4.48.0.3 or 4.16.0.2, while if $p=3$, then $E$ must come from a rational point on 9.27.0.1.
Remark 1.4.
We stop for a moment and point out that elliptic curves corresponding to rational points on 4.16.0.2 have the property that $\mathbb{Q}(E[4])=\mathbb{Q}(E[2])=\mathbb{Q}(E[2],i).$ Thus, these curves actually have a coincidence of division fields. In contrast, the elliptic curves corresponding to rational points on 4.48.0.3 have the property that $\mathbb{Q}(E[2])=\mathbb{Q}$ and $\mathbb{Q}(E[4])=\mathbb{Q}(i)$. So these curves have a near coincidence without having a coincidence of division fields. It is interesting to note that in order to have a near coincidence between the $2$- and $4$-division fields, without having an actual coincidence between the $2$- and $4$-division fields, the $2$-division field has to be trivial.
Similarly, the rational points on 9.27.0.1 yield elliptic curves with a near coincidence between their 3- and 9-division fields. Unlike the case with $p=2$, a rational point on 9.27.0.1 corresponds to an elliptic curve $E$ with either $j(E)=0$ or a surjective $\bar{\rho}_{E,3}$. In particular, if $E$ is a non-CM elliptic curve with $\mathbb{Q}(E[9])=\mathbb{Q}(E[3],\zeta_{9})$, we must have $[\mathbb{Q}(E[3]):\mathbb{Q}]=48$.
Theorem 1.3 ends up being the key ingredient in the answer of a seemingly unrelated question.
@@
Let $K/k$ be a Galois extension of fields. We say that $K/k$ is a nilpotent extension if $\operatorname{Gal}(K/k)$ is a nilpotent group. When the base field is obvious, we will just say that $K$ is a nilpotent field for brevity.
A classical result is that given $E/\mathbb{Q}$ and $n\in\mathbb{Z}^{+}$, the extension $\mathbb{Q}(E[n])/\mathbb{Q}$ is always a Galois extension. With this definition, a natural question one could ask is when is the extension $\mathbb{Q}(E[n])/\mathbb{Q}$ a nilpotent extension?
This question is not one that exists in a vacuum. Much effort has been spent studying extensions of the form $\mathbb{Q}(E[n])/\mathbb{Q}$. For example, in [14], González-Jiménez and Lozano-Robledo classify all elliptic curves $E/\mathbb{Q}$ such that $\mathbb{Q}(E[n])$ is contained in a cyclotomic extension of $\mathbb{Q}$ or, equivalently (by the Kronecker–Weber theorem), when $\mathbb{Q}(E[n])/\mathbb{Q}$ is an abelian extension.
We will be able to give a complete answer to the question of when $\mathbb{Q}(E[n])/\mathbb{Q}$ is nilpotent that is conditional on the following well-known conjecture.
@@
[37, Conjecture 1.1], [26, Conjecture 1.1.5] If $p>11$, then there is no non-CM elliptic curve $E/\mathbb{Q}$ for which the image of the mod $p$ Galois representation is contained in the normalizer of the non-split Cartan subgroup.
To state our next main result, recall that a Mersenne prime is a prime which is one less than a power of $2$. Such a prime has the form $12^{p}-$ 1 for $p$ a prime. A Fermat prime is a Fermat number (i.e. a number of the form $12^{2^{n}}+$ 1 with $n\geq 0$) which is prime.
@@
    If $E$ has complex multiplication by the order of discriminant $D=-27$, then $\mathbb{Q}(E[n])/\mathbb{Q}$ is never nilpotent.
@@
    $\begin{cases}d\equiv 1\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=2,\\ d\equiv 2\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=3,\\ d\equiv 2\cdot p^{\frac{p-1}{3}}\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=3\cdot 2^{k}+1\hbox{ for some $k\geq 1$},\\ d\equiv 2\cdot p^{\frac{p+1}{3}}\pmod{(\mathbb{Q}^{\times})^{3}}\text{ if }p=3\cdot 2^{k}-1\hbox{ for some $k\geq 1$}.\\ \end{cases}$
The reason that Conjecture 1.6 is needed is that if $p$ is a Mersenne prime, then the normalizer of the non-split Cartan subgroup in $\operatorname{GL}_{2}(\mathbb{F}_{p})$ is nilpotent. Theorem 1.7 part (2) is the strongest result we are able to prove without the assumption of Conjecture 1.6.
A consequence of Theorem 1.7 is the following result.
@@
There is no elliptic curve $E/\mathbb{Q}$ such that $\mathbb{Q}(E[19])/\mathbb{Q}$ is a nilpotent extension. Further, 19 is the smallest prime with this property.
@@
Our analysis of when $\mathbb{Q}(E[n])/\mathbb{Q}$ is nilpotent yields the following observation. If $E\colon y^{2}=x^{3}-x$ is the congruent number elliptic curve, then the points in $E[n]-\{O\}$ are constructible (i.e. obtainable from 00 and $1$ in finitely many steps using a compass and an unmarked straightedge) if and only if $n$ is a power of $2$ times a product of distinct Fermat primes. Since $\mathbb{Q}(\zeta_{n})\subseteq\mathbb{Q}(E[n])$, this is an extension of the Gauss-Wantzel theorem that a regular $n$-gon is constructible if and only if $n$ is a power of two times a product of distinct Fermat primes.
@@ 1.1. Outline of the paper
In Section 2 we remind the readers about the basic facts related to elliptic and modular curves that will be necessary for the proof of the main results. The proof of Theorem 1.3 will be handled in Section 3 by finding the smallest power $n$ such that $\mathbb{Q}(E[p^{n+1}])$ cannot be $\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}})$ and then we prove that if $\mathbb{Q}(E[p^{n+1}])\neq\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}})$, then $\mathbb{Q}(E[p^{n+2}])\neq\mathbb{Q}(E[p^{n+1}],\zeta_{p^{n+2}})$. The proof will have to be broken down into cases depending on if $p=2,3,5,$ or $p\geq 7.$
The proof of Theorem 1.7 starts in Section 4 by using group theory to classify the nilpotent subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ according to their image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$. In the same section we then apply what is known about the corresponding modular curves to say as much as we can about when $\mathbb{Q}(E[p])/\mathbb{Q}$ can be a nilpotent extension. This is where we will employ Conjecture 1.6.
@@
All of the computations in this paper were performed using [4] and the code can be found at [8].

### v3

@@ Abstract.
Let $E/\mathbb{Q}$ be an elliptic curve. We say that $E$ has a near coincidence of level $(n,m)$ if $m\mid n$ and $\mathbb{Q}(E[n])=\mathbb{Q}(E[m],\zeta_{n})$. We classify near coincidences of prime power level and use this result to give a classification of values of $n$ for which $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ is a nilpotent group. Along the way we prove a Gauss-Wantzel analog for the elliptic curve $E\colon y^{2}=x^{3}-x$, showing that $\mathbb{Q}(E[n])/\mathbb{Q}$ is constructible if and only if $\varphi(n)$ is a power of 2. Assuming that there are no non-CM rational points on the modular curves $X_{ns}^{+}(p)$ for primes $p>11$, we show that $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ nilpotent implies that $n$ is a power of $2$ or $n\in\{3,5,6,7,15,21\}$.
@@ 1. Introduction
For much of mathematical history, compass and straightedge constructions have served as exemplars of mathematical reasoning. At the center of the study of compass and straightedge constructions is the question of which regular polygons are constructible. First considered by the Greeks, an answer to this question remained elusive to mathematicians for hundreds of years. For centuries after this question was first considered, no progress was made. This changed when Gauss proved the following theorem:
@@
[15] Suppose that $n$ is of the form
$n=2^{k}\cdot p_{1}\cdots p_{j}$
where $p_{i}$ is a Fermat prime. Then it is possible to construct a regular $n$-gon using only a straightedge and compass.
Here we remind the reader that a Fermat prime is a Fermat number (i.e., a number of the form $2^{2^{m}}+1$ with $m\geq 0$) which is prime.
In fact, in [15] Gauss states that he has a proof of the converse, but he says, “the limits of the present work exclude this demonstration here.” The final part of the question was formally answered in 1837 when Pierre Wantzel proved the converse.
@@
[43] Suppose $n$ is a positive integer such that a regular $n$-gon is constructible. Then it must be that $n$ is of the form
$n=2^{k}\cdot p_{1}\cdots p_{j}$
where each $p_{i}$ is a Fermat prime.
Together, these two results together form the Gauss–Wantzel theorem. If we let $\varphi$ be Euler’s totient function, then we can summarize the Gauss–Wantzel theorem as saying a regular $n$-gon is constructible if and only if there exists a $k$ such that $\varphi(n)=2^{k}$.
A modern reformulation of this is to take the standard plane and turn it into the complex plane. A point in the complex plane is said to be constructible if there is a way to define that point using just a straightedge and compass. Interpreting the Gauss–Wantzel theorem through this more modern lens gives us a new interpretation of what it means for a complex number to be constructible.
Theorem 1.3.
[32, Theorem 4.53] Let $\alpha\in\mathbb{C}$. Then, $\alpha$ is constructible if and only if there exists an ascending chain of fields
$\mathbb{Q}=K_{0}\subseteq K_{1}\subseteq K_{2}\subset\dots\subseteq K_{j}$
such that $\alpha\in K_{j}$ and $K_{i}\colon K_{i-1}]=2.$
A subfield of $\mathbb{C}$ is said to be constructible if all of the elements in the field are constructible. Using Galois theory from here we can update Theorem 1.3 to say that $\alpha$ is constructible if and only if there is a Galois extension $K/\mathbb{Q}$ that contains $\alpha$ such that $\operatorname{Gal}(K/\mathbb{Q})$ is a 2-group. Given that the $n$th roots of unity are equally spaced around the unit circle, from this updated perspective, the question of about the constructability of a regular $n$-gon is exactly the same as the question about the constructability of $\mathbb{Q}(\zeta_{n})$. Here $\zeta_{n}$ is a primitive $n$th root of unity. This modern perspective allows us to examine the constructability of more complicated fields. The primary objects of study in this paper are the division fields of elliptic curves.
First we fix $\overline{\mathbb{Q}}$, an algebraic closure of $\mathbb{Q}$ and let $E/\mathbb{Q}$ be an elliptic curve. A classical result is that the points on $E$ can be given the structure of an abelian group. We let $E[n]$ be the $n$-torsion on $E$ defined over $\overline{\mathbb{Q}}$, that is $E[n]=\{P\in E(\overline{Q})\colon n\cdot P=\mathcal{O}\}$. Then we can define $\mathbb{Q}(E[n])$ to be the field of definition of the $n$-torsion points. With this we present the following theorem.
Theorem 1.4.
(A Gauss-Wantzel Analog) Let $E$ be the elliptic curve given by $y^{2}=x^{3}-x$ and let $n\geq 2$. Then, $\mathbb{Q}(E[n])$ is a constructible field if and only if $\varphi(n)=2^{k}$ for some integer $k$.
Proof.
Suppose $n\in\mathbb{Z}$ such that $\mathbb{Q}(E[n])/\mathbb{Q}$ is a constructible extension. By properties of the Weil pairing, $\mathbb{Q}(\zeta_{n})\subseteq\mathbb{Q}(E[n])$, and so by the Gauss–Wantzel Theorem it must be that $\varphi(n)$ is a power of 2.
In the opposite direction, suppose that $n\in\mathbb{Z}$ such that $\varphi(n)$ is a power of 2. Given that constructability is closed under compositum, it is enough to understand when $\mathbb{Q}(E[p^{k}])$ is constructible for some prime power $p^{k}$.
Suppose $p=2$. In this case, $\mathbb{Q}(E[2])=\mathbb{Q}$ and $\mathbb{Q}(E[2^{k+1}]):\mathbb{Q}(E[2^{k}])]=2^{m}$ for some $m\in\mathbb{Z}$ depending on $k$. Thus, for any $k$, $\operatorname{Gal}(\mathbb{Q}(E[2^{k}])/\mathbb{Q})$ is a 2-group and hence constructible.
Next suppose that $p=2^{2^{n}}+1$ is a Fermat prime and $E:y^{2}=x^{3}-x$. We start this case by showing that $\mathbb{Q}(E[p],i):\mathbb{Q}(i)]$ is a power of $2$, which implies the desired result. The endomorphism ring of $E$ over $\mathbb{Q}(i)$ is isomorphic to $\mathbb{Z}[i]$. Moreover, since $p\equiv 1\pmod{4}$ splits we can factor $p=\pi\overline{\pi}$ for $\pi\in\mathbb{Z}[i]$. (Explicitly, we have $\pi=2^{2^{n-1}}+i$ and $\overline{\pi}=2^{2^{n-1}}-i$.) We therefore have two isogenies $\phi_{1}\colon E\to E$ and $\phi_{2}\colon E\to E$ given by $\phi_{1}(P)=[\pi]P$ and $\phi_{2}(P)=[\overline{\pi}]P$ which are defined over $\mathbb{Q}(i)$. Let $E[\pi]$ and $E[\overline{\pi}]$ be the kernels of these two isogenies. It is straightforward to see that $E[p]$ is the direct product of $E[\pi]$ and $E[\overline{\pi}]$, that each of $E[\pi]$ and $E[\overline{\pi}]$ has order $p$, and both $E[\pi]$ and $E[\overline{\pi}]$ are Galois stable over $\mathbb{Q}(i)$. This shows that the image of the mod $p$ Galois representation $\rho_{E,p}\colon\operatorname{Gal}(\mathbb{Q}(E[p],i)/\mathbb{Q}(i))\to\operatorname{GL}_{2}(\mathbb{F}_{p})$ is conjugate to a subgroup of the diagonal matrices. The set of diagonal matrices has order $(p-1)^{2}=2^{2^{n+1}}$. Since $\rho$ is injective, Lagrange’s theorem implies that $\mathbb{Q}(E[p],i):\mathbb{Q}(i)]$ is a power of $2$. ∎
One might think that the more general question of when $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ can be a $p$-group would add some interest, but it turns out the only interesting case here is when $p=2$. This is easy to see since $\mathbb{Q}(E[n]):\mathbb{Q}]$ is even for all $n\geq 3$. If we want to study a condition on $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ that is more general than being a $p$-group, but is restrictive enough to yield meaningful results, a naive thing to do would be to study when $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ is the direct product of a finite number of $p$-groups. This allows some additional flexibility, but still imposes some structure on the situation.
While at first this seems arbitrary and naive, this is a very natural thing to do.
Proposition 1.5.
[18, Proposition II.7.5] Let $G$ be a finite group, then $G$ is nilpotent if and only if it is the direct product of its $p$-Sylow subgroups.
Thus, our naive consideration is actually logically equivalent to the condition that $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ is nilpotent. Looking at the literature, we find the following result about when $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ is abelian.
Theorem 1.6.
[16, Theorem 1.1] Let $E/\mathbb{Q}$ be an elliptic curve and $n\geq 2$. If $\mathbb{Q}(E[n])/\mathbb{Q}$ is an abelian extension, then $n=2,3,4,5,6,$ or $8$.
With this added context, our new goal is to understand when $\operatorname{Gal}(\mathbb{Q}(E[n])/\mathbb{Q})$ is nilpotent for $E/\mathbb{Q}$. To that end, we give the following definition:
@@
Let $K/k$ be a Galois extension of fields. We say that $K/k$ is a nilpotent extension if $\operatorname{Gal}({K}/k)$ is a nilpotent group. When the base field $k$ is clear from context, we will just say that $K$ is a nilpotent field for brevity.
We will be able to give a complete answer to the question of when $\mathbb{Q}(E[n])/\mathbb{Q}$ is nilpotent that is conditional on a positive answer to Serre’s uniformity question.
@@
[46, Conjecture 1.1], [33, Conjecture 1.1.5] If $p>11$, then there is no non-CM elliptic curve $E/\mathbb{Q}$ for which the image of the mod $p$ Galois representation is contained in the normalizer of the non-split Cartan subgroup.
Before stating our next theorem we remind the reader that a Mersenne prime is a prime which is one less than a power of $2$. Such a prime always has the form $2^{p}-1$ for $p$ a prime.
@@
    If $E$ has complex multiplication by the order of discriminant D=27D=-27, then $\mathbb{Q}(E[n])/\mathbb{Q}$ is never nilpotent.
@@
[…]
One corollary of this result is the following.
@@
There is no elliptic curve $E/\mathbb{Q}$ for which $\mathbb{Q}(E[19])/\mathbb{Q}$ is a nilpotent extension. Further, $19$ is the smallest prime with this property.
Along the way to proving Theorem 1.9, a phenomenon appears that we point out here. When studying the question of when $\mathbb{Q}(E[p^{n}])/\mathbb{Q}$ nilpotent given that $\mathbb{Q}(E[p^{n-1}])/\mathbb{Q}$ is, we realized that we would need to classify exactly when $\mathbb{Q}(E[p^{n}])=\mathbb{Q}(E[p^{n-1}],\zeta_{p^{n}})$.
Definition 1.11.
Let $E/\mathbb{Q}$ be an elliptic curve and let $m,n\in\mathbb{Z}^{+}$ such that $m\mid n$. We say that there is a near coincidence between the $m$- and $n$-division fields of $E$ if
$\mathbb{Q}(E[n])=\mathbb{Q}(E[m],\zeta_{n}).$
The terminology here is motivated by the definition given in [7] where an elliptic curve $E/\mathbb{Q}$ is said to have a coincidence between its division fields if $\mathbb{Q}(E[m])=\mathbb{Q}(E[n])$ for distinct integers $m$ and $n$. In that paper is the following theorem:
Theorem 1.12.
[7, Theorem 1.4] Let $E/\mathbb{Q}$ be an elliptic curve, $p$ be a prime, and let $n\in\mathbb{Z}^{+}.$
  1. (1)
    Suppose $\mathbb{Q}(E[p^{n+1}])=\mathbb{Q}(E[p^{n}]).$ Then $p=2$ and $n=1.$
  2. (2)
The problem of classifying such coincidences for elliptic curves defined over number fields was also recently investigated by Yvon in [44].
In the end, we are able to give a complete classification of elliptic curves over $\mathbb{Q}$ with a near coincidence of their $p^{n+1}$ and $p^{n}$ division fields; i.e., when $\mathbb{Q}(E[p^{n+1}])=\mathbb{Q}(E[p^{n}],\zeta_{p}^{n+1}).$
Theorem 1.13.
$\mathbb{Q}(E[p^{n+1}])=\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}}).$
Then $p\in\{2,3\}$ and $n=1$. Further, if $p=2$, then $E$ must correspond to a rational point on the one of the modular curves with RSZB label 4.48.0.3 or 4.16.0.2, while if $p=3$, then $E$ must come from a rational point on 9.27.0.1.
@@
Before moving on, we point out that elliptic curves corresponding to rational points on 4.16.0.2 have the property that $\mathbb{Q}(E[4])=\mathbb{Q}(E[2])=\mathbb{Q}(E[2],i).$ Thus, these curves actually have a coincidence of division fields. In contrast, the elliptic curves corresponding to rational points on 4.48.0.3 have the property that $\mathbb{Q}(E[2])=\mathbb{Q}$ and $\mathbb{Q}(E[4])=\mathbb{Q}(i)$. So these curves have a near coincidence without having a coincidence of division fields. It is interesting to note that in order to have a near coincidence between the $2$- and $4$-division fields, without having an actual coincidence between the $2$- and $4$-division fields, the $2$-division field has to be trivial.
Similarly, the rational points on 9.27.0.1 yield elliptic curves with a near coincidence between their 3- and 9-division fields. Unlike the case with $p=2$, a rational point on 9.27.0.1 corresponds to an elliptic curve $E$ with either $j(E)=0$ or a surjective $\bar{\rho}_{E,3}$. In particular, if $E$ is a non-CM elliptic curve with $\mathbb{Q}(E[9])=\mathbb{Q}(E[3],\zeta_{9})$, we must have $\mathbb{Q}(E[3]):\mathbb{Q}]=48$.
Remark 1.15.
We choose to begin this paper with the Gauss–Wantzel theorem not only because of its classical significance, but because it serves as a natural entry point for a broad audience into the kinds of questions that we hope to address in this paper. The main objective here is to classify when the division field $\mathbb{Q}(E[n])/\mathbb{Q}$ of an elliptic curve $E$ over $\mathbb{Q}$ is nilpotent. Along the way, we uncover a surprising analog of the Gauss–Wantzel theorem as well as a result about near coincidences of division fields. We hope that the Gauss–Wantzel analog, while not the main thrust of the paper, illustrates the value of considering questions like these.
@@ 1.1. Outline of the paper
In Section 2 we recall basic facts related to elliptic and modular curves that will be necessary for the proof of the main results. The proof of Theorem 1.13 will be handled in Section 3 by finding the smallest power $n$ such that $\mathbb{Q}(E[p^{n+1}])$ cannot be $\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}})$ and then we prove that if $\mathbb{Q}(E[p^{n+1}])\neq\mathbb{Q}(E[p^{n}],\zeta_{p^{n+1}})$, then $\mathbb{Q}(E[p^{n+2}])\neq\mathbb{Q}(E[p^{n+1}],\zeta_{p^{n+2}})$. The proof will have to be broken down into cases depending on if $p=2,3,5,$ or $p\geq 7.$
The proof of Theorem 1.9 starts in Section 4 by using group theory to classify the nilpotent subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/p\mathbb{Z})$ according to their image in $\operatorname{PGL}_{2}(\mathbb{Z}/p\mathbb{Z})$. We then apply known results about the corresponding modular curves to determine when the extension $\mathbb{Q}(E[p])/\mathbb{Q}$ can be nilpotent. This is where we will employ Conjecture 1.8.
@@
All of the computations in this paper were performed using Magma [4] and the code can be found at [8].